C-2 Manufacturing The capacity in gallons of a cylindrical tank can be found using the formula
C=pieD^2L/924
where
C=capacity in gallons
D=diameter of tank in inches
L=length of tank in inches
(a) Find the diameter of a 42-in-long tank that has a capacity of 30 gallons.
(b) Find the diameter of a 60-in-long tank that has a capacity of 50 gallons.

Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

30=3.14d^2*42/924
30=131.88d^2/924 now cross multiply
30*924=131.88d^2
27,720=131.88d^2 now divide both sides by 131.88
d^2=27,720/131.88
d^2=210.191
d=sqrt210.191
d=14.5 in. is the diameter.
-----------------------------------------
50=3.14d^2*60/924
50=188.4d^2/924
188.4d^2=50*924
188.4d^2=46,200
d^2=46,200/188.4
d^2=245.22
d=sqrt245.22
d=15.66 in. is the diameter.

Answer by stanbon(75887)

C=pieD^2L/924
where
C=capacity in gallons
D=diameter of tank in inches
L=length of tank in inches
(a) Find the diameter of a 42-in-long tank that has a capacity of 30 gallons.
(b) Find the diameter of a 60-in-long tank that has a capacity of 50 gallons.
--------
C=pieD^2L/924
Solve for D:
D = sqrt[924C/(pi)L]
------
a) D = sqrt[924*30/[(pi)*42]
D = 21.08 inches
------------
b) D = sqrt[924*50/[pi*60]
D = 245.0986 inches
==============
Cheers,
Stan H.